# Mind the GAP! The Challenges of Scale in Pixel-based Deep Reinforcement Learning

### v1

@@ Abstract
Scaling deep reinforcement learning in pixel-based environments presents a significant challenge, often resulting in diminished performance. While recent works have proposed algorithmic and architectural approaches to address this, the underlying cause of the performance drop remains unclear. In this paper, we identify the connection between the output of the encoder (a stack of convolutional layers) and the ensuing dense layers as the main underlying factor limiting scaling capabilities; we denote this connection as the bottleneck, and we demonstrate that previous approaches implicitly target this bottleneck. As a result of our analyses, we present global average pooling as a simple yet effective way of targeting the bottleneck, thereby avoiding the complexity of earlier approaches.
@@ 1 Introduction
Reinforcement learning (RL) is widely considered one of the most effective approaches for complex sequential decision-making problems (Mnih et al., 2015; Vinyals et al., 2019; Bellemare et al., 2020; Degrave et al., 2022; Wurman et al., 2022), in particular when combined with deep neural networks (typically referred to as deep RL). In contrast to the so-called “scaling laws” observed in supervised learning (where larger networks typically result in improved performance), it is difficult to scale RL networks without sacrificing performance. There has been a recent line of work aimed at developing techniques for effectively scaling value-based networks, such as via the use of mixtures-of-experts (Obando Ceron* et al., 2024), network pruning (Obando Ceron et al., 2024), tokenization (Sokar et al., 2025), and regularization (Nauman et al., 2024). These techniques tend to focus on structural modifications to standard deep RL networks by leveraging sparse-network training techniques from the supervised learning literature.
@@
One unifying aspect of the aforementioned works is that they tend to be most effective on networks that process pixel inputs, such as when training on the ALE (Bellemare et al., 2013). These networks are typically divided into an encoder $\phi$ consisting of a series of convolutional layers, followed by a series of dense layers $\psi$; thus, for an input $x$, the network output is given by $\psi(\phi(x))$. Obando Ceron* et al. (2024) and Sokar et al. (2025) scaled the first layer of $\psi$ while Obando Ceron et al. (2024) scaled all layers in $\phi$. It is worth highlighting that $\psi\circ\phi$ is, in practice, a set of weights connecting the output of $\phi(x)$ with the input layer of $\psi$.
@@
Our contributions can be summarized as follows: (i) We investigate the challenges leading to the performance degradation of scaled RL networks; (ii) we study the underlying reasons behind the success of existing architectural approaches in scaling; and (iii) we present pooling as a faster, simpler method yielding superior performance. We begin in Section 2 by providing the necessary background on Reinforcement Learning (RL), detailing the architecture employed in pixel-based deep RL, and outlining architectural modifications proposed by recent methods for addressing scaling challenges. Section 3.1 then describes our experimental setup. Our analyses investigating the difficulties of scaling RL networks and how existing methods attempt to address them are presented in Sections 3.2 and 3.3, respectively. Section 4 is dedicated to our results and analysis on RL networks with GAP. Finally, Section 5 covers the related work, followed by a conclusion and discussion in Section 6.
@@ 2.1 Reinforcement learning
Reinforcement learning (RL) involves an agent moving through a series of states $x\in\mathcal{X}$ by selecting an action $a\in\mathcal{A}$ at discrete timesteps. After selecting action $a_{t}$ from state $x_{t}$, the agent will receive a reward $r_{t}(x_{t},a_{t})$, and its goal is to maximize the discounted sum of cumulative rewards $\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma\in[0,1)$ by finding an optimal policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ which quantifies the agent’s behavior at each state. Value-based methods (Sutton and Barto, 1998) maintain estimates of the value of selecting action $a$ from state $x$ and following $\pi$ afterwards: $Q(x,a):=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}(x_{t},a_{t})|x_{0}=x,a_{0}=a,a_{t}\sim\pi(x_{t})\right]$, where $\pi$ is induced from $Q$, for instance with the use of softmax: $\pi(x)(a):=\frac{e^{Q(x,a)}}{\sum_{a^{\prime}\in\mathcal{A}}e^{Q(x,a^{\prime})}}$.
Mnih et al. (2015) demonstrated deep neural networks can be very effective at approximating $Q$-values, even for complex domains such as Atari games (Bellemare et al., 2013); their network has served as the backbone for most deep RL network. Later, Espeholt et al. (2018) proposed using ResNet based architecture which demonstrates significant performance improvements over the original CNN architecture. For pixel-based environments, this family of networks consists of a set of convolutional layers (Fukushima, 1980), which we will collectively refer to as $\phi$ (and often referred to as the encoder or representation), followed by a set of dense layers, which we will collectively refer to as $\psi$. Thus given an input $x$, the network approximates the $Q$-values as $\tilde{Q}(x,\cdot)=\psi(\phi(x))$.
@@
2.2 SoftMoE
Obando Ceron* et al. (2024) proposed replacing the dense layer in $\psi$ with mixture-of-experts (MoEs). The input to the mixture is a set of tokens created by restructuring the 3-dimensional output of $\phi$. Their investigation into different tokenization strategies revealed that forming $H\times W$ tokens of $C$ dimensions achieved the highest performance. This reduces the number of parameters of the bottleneck to $C\times dim(\psi)$. Sokar et al. (2025) demonstrated that tokenzing the output of $\psi$ is a core source of the performance improvements of MoEs in RL and single expert (SoftMoE-1) has a close performance to using multiple experts across different scales.
2.3 Sparse Methods
Sokar et al. (2022) demonstrated that using sparse neural networks instead of dense networks increases learning speed and leads to improved performance. Graesser et al. (2022) studied different ways to induce sparsity in the network including pruning the network weights during training and using sparse networks from scratch. With a sparsity level $s$, the number of parameters in the bottleneck is $s\times H\times W\times C\times dim(\psi)$. Recently (Obando Ceron et al., 2024) demonstrated that pruning is an effective approach that enables scaling networks. Below we will explain how these approaches change the architecture of the network.
@@ 2.3 Sparse Methods
The training starts with a dense network and the low magnitude parameters are progressively pruned throughout training under a polynomial schedule (Zhu and Gupta, 2017). At the end iteration of pruning, the network will reach the target sparsity which is kept fixed during the rest of training.
@@
A sparse neural network is defined at initialization with specific sparsity level that is maintained throughout training. The sparse topology is either fixed known as static, or dynamically optimized during training through a periodic change of a portion of the connections as proposed in RigL (Evci et al., 2020).
@@ 3 Analyses
Main hypothesis A low-dimensional and well-structured bottleneck enables scaling deep RL networks.
@@ 3.1 Experimental setup
We use the Impala architecture (Espeholt et al., 2018), as it has been demonstrated to yield stronger performance than the original architecture of Mnih et al. (2015). Our main experiments and analyses show results across different scales for the width of $\psi$ (×$\times 1$, ×$\times 2$, ×$\times 4$, ×$\times 8$). When a scale is not explicitly stated, our analysis defaults to the $\times$4 scale due to its observed high gain (Obando Ceron* et al., 2024; Sokar et al., 2025; Obando Ceron et al., 2024). Following (Obando Ceron* et al., 2024; Sokar et al., 2025), we only scale $\psi$, unless specified otherwise. While our main focus is on the more performant ResNet backbone of Impala, we also demonstrate the broader applicability of our findings using a standard CNN architecture (Mnih et al., 2015).
@@
We use the Rainbow agent (Hessel et al., 2018) and evaluate on the Arcade Learning Environment (ALE) suite (Bellemare et al., 2013) using the same $20$ Atari games used by Obando Ceron* et al. (2024) and Sokar et al. (2025). Moreover, we present our main results on the full suite of 60 games. We ran each experiment for a total of 200 million environment steps, with $5$ independent seeds, except for the experiments with an increased replay ratio of 2, where we report the results at 50M steps. Moreover, we evaluate Rainbow on the Procgen (Cobbe et al., 2019) benchmark and on the 100k benchmark (Łukasz Kaiser et al., 2020) using DER (the version of Rainbow tuned for this setting).
@@
Figure 2: (Left) Distribution of dormant neurons across $\phi$ and $\psi$ in scaled baseline across different games at the end of training. The fully connected layer exhibits the highest percentage of dormancy. (Right) The performance degradation associated with scaling the entire network architecture is similar to that observed when only the bottleneck is scaled.
@@
For all algorithms, we use the default hyperparameters in the Dopamine library. For sparse-training algorithms, we follow (Graesser et al., 2022) and use 90% sparsity. For gradual pruning, we start pruning at 8M environment steps and stop at 160M (80% into training), following the recommendation of Graesser et al. (2022); Obando Ceron et al. (2024). For RigL (Evci et al., 2020), we use a drop fraction of 20% and an update interval of 5000. For SoftMoE, we use SoftMoE-1 not only because it has been shown to have a similar performance as standard SoftMoE (Sokar et al., 2025), but also to ensure we have a similar architecture as the other methods considered. We report interquartile mean (IQM) with 95% stratified bootstrap confidence intervals as recommended in (Agarwal et al., 2021). Full experimental details are provided in the appendix.
@@ 3.2 Why scaling deep RL networks hurts performance
We start by analyzing the training dynamics of a network where all layers within $\phi$ and $\psi$ are scaled by a factor of $4$. We examine neuron activity by measuring the fraction of dormant neurons (Sokar et al., 2023), which is an indicator of plasticity. We find that $\psi$ exhibits a large fraction of dormant neurons, while $\phi$ has low dormancy rates, as shown in the left plot of Figure 2. This suggests that scaling mostly affects the plasticity of the bottleneck.
In the right plot of Figure 2, we compare the performance of the baseline network against the same network with scaled $\phi$ and $\psi$. As consistently observed in previous studies, the scaled network exhibits a degradation in performance. To investigate the contribution of $\psi$ in this performance decrease, we only scale the bottleneck (via the first layer of $\psi$) to match the parameter count of the fully scaled network. As can be observed, the performance of the two scaled models is nearly identical, suggesting that the bottleneck drives most of the performance degradation when scaling.
@@
To further investigate potential challenges in feature learning within $\psi$, we study the network behavior when providing higher-level features to $\psi$. Specifically, we increased the depth of $\phi$ by adding four additional ResNet blocks, while using the scaled bottleneck. This modification makes $\psi$ receive more structured, high-level features from the encoder. We present the performance throughout training in Figure 4. The fact that we observe a dramatic increase in performance suggests that structured representations helps feature learning in scaled networks.
@@ 3.3 Existing techniques are mainly targetting the bottleneck
To validate our hypothesis on sparse methods, we evaluated the impact of sparse training techniques when limited to the bottleneck, while keeping all other layers dense. We performed this analysis on gradual pruning (Obando Ceron et al., 2024), dynamic sparsity (RigL) and static sparsity (Graesser et al., 2022; Sokar et al., 2022). As shown in the left plot of Figure 5, restricting sparsification to the scaled bottleneck results in improved performance across all sparse training techniques. This suggests that applying sparsity only to the bottleneck is sufficient to enable scaling RL networks.
Figure 5: (Left) Across different sparse algorithms, sparsification of only $\psi$ yields better performance than sparsifying $\phi$ and $\psi$. (Right) The relation between performance and the effective number of parameters in $\psi$ for different approaches. Architectural methods have lower effective density than the baseline which correlates with the observed performance improvements.
In addition to structuring the encoder output, existing techniques effectively reduce the dimensionality of the bottleneck by either structuring the output as tokens (as in SoftMoE) which results in a lower dimensional input to $\psi$, or explicitly masking the majority of input weights (as in sparse methods). The right plot of Figure 5 confirms this reduction in parameters, and illustrates a positive correlation between scale and performance, which is notably absent in the baseline. This suggests that low-dimensional and structured bottlenecks facilitate scaling deep RL networks.

### v2

@@ Abstract
Scaling deep reinforcement learning in pixel-based environments presents a significant challenge, often resulting in diminished performance. While recent works have proposed algorithmic and architectural approaches to address this, the underlying cause of the performance drop remains unclear. In this paper, we identify the connection between the output of the encoder (a stack of convolutional layers) and the ensuing dense layers as the main underlying factor limiting scaling capabilities; we denote this connection as the bottleneck, and we demonstrate that previous approaches implicitly target this bottleneck. As a result of our analyses, we present Global Average Pooling (GAP) as a simple yet effective way of targeting the bottleneck, thereby avoiding the complexity of earlier approaches.
@@ 1 Introduction
Reinforcement Learning (RL) is widely considered one of the most effective approaches for complex sequential decision-making problems (Mnih et al., 2015; Vinyals et al., 2019; Bellemare et al., 2020; Degrave et al., 2022; Wurman et al., 2022), in particular when combined with deep neural networks (typically referred to as deep RL). In contrast to the so-called “scaling laws” observed in supervised learning (where larger networks typically result in improved performance) (Kaplan et al., 2020), it is difficult to scale RL networks without sacrificing performance. There has been a recent line of work aimed at developing techniques for effectively scaling value-based networks, such as via the use of mixtures-of-experts (Obando Ceron* et al., 2024), network pruning (Obando Ceron et al., 2024), tokenization (Sokar et al., 2025), and regularization (Nauman et al., 2024). Most of these techniques tend to focus on structural modifications to standard deep RL networks by leveraging sparse-network training techniques from the supervised learning literature.
@@
One unifying aspect of the aforementioned works is that they tend to be most effective on networks that process pixel inputs, such as when training on the Arcade Learning Environment (ALE) (Bellemare et al., 2013). These networks are typically divided into an encoder $\phi$ consisting of a series of convolutional layers, followed by a series of dense layers $\psi$; thus, for an input $x$, the network output is given by $\psi(\phi(x))$. Obando Ceron* et al. (2024) and Sokar et al. (2025) scaled the first layer of $\psi$ while Obando Ceron et al. (2024) scaled all layers in $\phi$. It is worth highlighting that $\psi\circ\phi$ is, in practice, a set of weights connecting the output of $\phi(x)$ with the input layer of $\psi$.
@@
Our contributions can be summarized as follows: (i) We investigate the challenges leading to the performance degradation of scaled RL networks; (ii) we study the underlying reasons behind the success of existing architectural approaches in scaling; and (iii) we present pooling as a faster, simpler method yielding superior performance. We begin in Section 2 by providing the necessary background on reinforcement learning, detailing the architecture employed in pixel-based deep RL, and outlining architectural modifications proposed by recent methods for addressing scaling challenges. Section 3.1 then describes our experimental setup. Our analyses investigating the difficulties of scaling RL networks and how existing methods attempt to address them are presented in Sections 3.2 and 3.3, respectively. Section 4 is dedicated to our results and analysis on RL networks with GAP. Finally, Section 5 covers the related work, followed by a conclusion and discussion in Section 6.
@@ 2.1 Reinforcement learning
Reinforcement learning involves an agent moving through a series of states $x\in\mathcal{X}$ by selecting an action $a\in\mathcal{A}$ at discrete timesteps. After selecting action $a_{t}$ from state $x_{t}$, the agent will receive a reward $r_{t}(x_{t},a_{t})$, and its goal is to maximize the discounted sum of cumulative rewards $\sum_{t=0}^{\infty}\gamma^{t}r_{t}$, where $\gamma\in[0,1)$ by finding an optimal policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{A})$ which quantifies the agent’s behavior at each state. Value-based methods (Sutton and Barto, 1998) maintain estimates of the value of selecting action $a$ from state $x$ and following $\pi$ afterwards: $Q(x,a):=\mathbb{E}\left[\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}(x_{t},a_{t})|x_{0}=x,a_{0}=a,a_{t}\sim\pi(x_{t})\right]\right]$, where $\pi$ is induced from $Q$, for instance with the use of softmax: $\pi(x)(a):=\frac{e^{Q(x,a)}}{\sum_{a^{\prime}\in\mathcal{A}}e^{Q(x,a^{\prime})}}$.
Mnih et al. (2015) demonstrated deep neural networks can be very effective at approximating $Q$-values, even for complex domains such as Atari games (Bellemare et al., 2013); their network has served as the backbone for most deep RL networks. Later, Espeholt et al. (2018) proposed using ResNet based architecture which demonstrates significant performance improvements over the original convolutional neural network (CNN) architecture. For pixel-based environments, this family of networks consists of a set of convolutional layers (Fukushima, 1980), which we will collectively refer to as $\phi$ (and often referred to as the encoder or representation), followed by a set of dense layers, which we will collectively refer to as $\psi$. Thus given an input $x$, the network approximates the $Q$-values as $\tilde{Q}(x,\cdot)=\psi(\phi(x))$.
@@
2.2 Network scaling
As we will demonstrate below, this bottleneck has a direct impact on learning efficiency. The standard approach is to flatten the output of $\phi$ into a single vector before feeding it to $\psi$, which we refer to as an ‘unstructured’ representation. This unstructured representation risks diluting the spatial structure captured by $\phi$, making learning more difficult. This is exacerbated when scaling the width of $\psi$, as demonstrated in prior works (Obando Ceron* et al., 2024; Sokar et al., 2025). We hypothesize that the effectiveness of prior architectural modifications for scaling RL networks lies in their ability to induce some form of structure at this bottleneck. The rest of this section details these specific modifications.
2.2.1 SoftMoE
To address scaling limitations, Obando Ceron* et al. (2024) proposed a notable architectural change: replacing the standard dense layer in $\psi$ with a mixture-of-experts (MoEs). This required first restructuring the 3-dimensional output of the encoder ($\phi$) into a set of tokens to be fed into the mixture. Their investigation into tokenization strategies concluded that forming $H\times W$ tokens, each with $C$ dimensions (the channel depth), yielded the best performance. This restructuring effectively changes the connection, reducing the bottleneck’s parameter count to $C\times dim(\psi)$ (See Figure 1). This architectural change has proven highly effective for various agents and at different scales.
Building on this, Sokar et al. (2025) isolated the source of these performance gains. They demonstrated that the tokenization step itself—not the MoE architecture—was the most critical component. Their evidence showed that a single expert model (SoftMoE-1), which retains the tokenization but omits the expert routing, achieves performance nearly identical to that of the full multi-expert model across different scales.
2.2.2 Sparse networks
Sparse methods offer an alternative approach to managing network complexity. Sokar et al. (2022) demonstrated that using sparse neural networks in place of dense ones can increase learning speed and lead to improved performance. Following this, Graesser et al. (2022) studied various ways to induce this sparsity, such as by pruning network weights during training or by using networks that are sparse from scratch. By imposing sparsity, these techniques directly structure the bottleneck, reducing the density of its connections. With a given sparsity level $s$, the effective number of parameters in the bottleneck is reduced to $s\times H\times W\times C\times dim(\psi)$. These sparse approaches have shown promise for the scaling of larger architectures; indeed, Obando Ceron et al. (2024) recently demonstrated that pruning is an effective approach that facilitates the scaling of networks.
@@ 2.2.2 Sparse networks
This strategy begins with a standard, fully-dense network. As training progresses, connections (parameters) with low magnitudes—which are considered less salient to the network’s function—are progressively removed. This pruning process often follows a polynomial schedule (Zhu and Gupta, 2017). Once the pruning schedule concludes, the network achieves its target sparsity level, and this fixed sparse architecture is maintained for the remainder of training.
@@
In contrast to gradual pruning, this approach defines a sparse network at initialization, and this specific sparsity level is maintained throughout training. The sparse topology can be static, meaning the set of active connections is fixed for the entire training duration. Alternatively, the topology can be dynamic. For example, the RigL method (Evci et al., 2020) dynamically optimizes the connections by periodically pruning a portion of the weights and growing new ones elsewhere, effectively "rewiring" the network while maintaining the overall sparsity.
@@ 3 Analyses
Main hypothesis A low-density and well-structured bottleneck enables scaling deep RL networks.
@@ 3.1 Experimental setup
We employ the Impala architecture (Espeholt et al., 2018) for its superior performance, and while our analysis centers on it, we also confirm our findings’ broad applicability using the standard CNN architecture (Mnih et al., 2015). Our main experiments and analyses present results across various scaling factors for the width of $\psi$ (specifically, $\times 1$, $\times 2$, $\times 4$, and $\times 8$). When a scaling factor is not explicitly mentioned, our analysis defaults to the $\times 4$ scale. This is because, as shown in prior work (Obando Ceron* et al., 2024; Sokar et al., 2025; Obando Ceron et al., 2024), the $\times 4$ scale provides a substantial performance improvement, while further scaling tends to yield only marginal gains. Although layer width is our primary focus, we also include some preliminary investigations into scaling network depth in the appendix.
@@
Our primary experimental setup involves the Rainbow agent (Hessel et al., 2018) evaluated on the Arcade Learning Environment (ALE) suite (Bellemare et al., 2013). For direct comparison with recent work, we use the same $20$-game subset from Obando Ceron* et al. (2024) and Sokar et al. (2025). However, we present our main results on the full ALE suite of 60 Atari games. We ran each experiment for a total of 200 million environment steps, with results averaged over $5$ independent seeds, except for the experiments with an increased replay ratio of 2, where we report the results at 50M steps. To further test the generality of our approach on discrete tasks, we also evaluate Rainbow on the Procgen benchmark (Cobbe et al., 2019). Additionally, we assess performance on the data-efficient 100k benchmark (Łukasz Kaiser et al., 2020), for which we use DER, a version of Rainbow specifically tuned for that data-constrained setting. Finally, to demonstrate our findings extend beyond discrete action spaces, we also evaluate the Soft Actor-Critic (SAC) agent (Haarnoja et al., 2018) on continuous control tasks from the DeepMind Control (DMC) suite (Tassa et al., 2018).
@@
Figure 2: (Left) Distribution of dormant neurons across $\phi$ and $\psi$ in scaled baseline across different games at the end of training. The fully connected layer exhibits the highest percentage of dormancy. (Right) The performance degradation associated with scaling the entire network architecture is comparable to that observed when only the bottleneck is scaled. The performance is aggregated over 20 games.
@@
For all algorithms, we use the default hyperparameters in the Dopamine library. For sparse-training algorithms, we follow (Graesser et al., 2022) and use 90% sparsity. Additional analysis covering other sparsity levels is included in the appendix. For gradual pruning, we start pruning at 8M environment steps and stop at 160M (80% into training), following the schedules recommended by Graesser et al. (2022) and Obando Ceron et al. (2024). For RigL (Evci et al., 2020), we use a drop fraction of 20% and a connection update interval of 5000. For SoftMoE, we use the single-expert variant (SoftMoE-1). This choice was twofold: first, it has been shown to yield performance comparable to the standard multi-expert SoftMoE (Sokar et al., 2025), and second, it ensures a more direct architectural comparison with the other methods considered in this study. We report interquartile mean (IQM) with 95% stratified bootstrap confidence intervals as recommended in (Agarwal et al., 2021). Full experimental details are provided in the appendix.
@@ 3.2 Why scaling deep RL networks hurts performance
We start by analyzing the training dynamics of a network where the width of all layers in both $\phi$ and $\psi$ are uniformly scaled by a factor of $4$. We examine neuron activity by measuring the fraction of dormant neurons, a metric that serves as a key indicator of network plasticity. Following Sokar et al. (2023), a neuron is considered "dormant" if its average activation falls below a certain threshold. As shown in the left plot of Figure 2, we find that $\psi$ exhibits a large fraction of dormant neurons, while $\phi$ has low dormancy rates. This suggests that scaling mostly affects the plasticity of the bottleneck.
In the right plot of Figure 2, we compare the performance of the baseline network against the same network with scaled $\phi$ and $\psi$. As consistently observed in previous studies, the scaled network exhibits a degradation in performance. To investigate the contribution of $\psi$ in this performance decrease, we only scale the bottleneck (via the first layer of $\psi$) to match the parameter count of the fully scaled network. As can be observed, the two scaled models has comparable performance, suggesting that the bottleneck drives most of the performance degradation when scaling.
@@
To further investigate potential challenges in feature learning within $\psi$, we study the network behavior when providing more abstract, higher-level features to $\psi$. Specifically, we increased the depth of $\phi$ by adding four additional ResNet blocks, while using the scaled bottleneck. This modification makes $\psi$ receive more structured, high-level features from the encoder. We present the performance throughout training in Figure 4. The fact that we observe a dramatic increase in performance when compared to the network that only scaled $\psi$ suggests that structured representations helps feature learning in scaled networks.
Figure 5: (Top) Across different sparse algorithms, sparsification of only $\psi$ yields better performance than sparsifying $\phi$ and $\psi$. (Bottom) The relation between performance and the effective number of parameters in $\psi$ for different approaches. Architectural methods have lower effective density than the baseline which correlates with the observed performance improvements.
@@ 3.3 Existing techniques are mainly targetting the bottleneck
To validate our hypothesis on sparse methods, we evaluated the impact of sparse training techniques when limited to the bottleneck, while keeping all other layers dense. We performed this analysis on gradual pruning (Obando Ceron et al., 2024), dynamic sparsity (RigL) and static sparsity (Graesser et al., 2022; Sokar et al., 2022). As shown in the top plot of Figure 5, restricting sparsification to the scaled bottleneck results in improved performance across all sparse training techniques. This suggests that applying sparsity only to the bottleneck is sufficient to enable scaling RL networks.
In addition to structuring the encoder output, existing techniques effectively reduce the density of the bottleneck by either structuring the output as tokens (as in SoftMoE) which results in a lower dimensional input to $\psi$, or explicitly masking the majority of input weights (as in sparse methods). The bottom plot of Figure 5 confirms this reduction in parameters, and illustrates a positive correlation between scale and performance, which is notably absent in the baseline. This suggests that low-density and structured bottlenecks facilitate scaling deep RL networks.